### v1

@@ Abstract
We analyze blocks proposed for inclusion in the Ethereum blockchain during 8 minutes on December 3rd, 2024. Our dataset comprises 38 winning blocks, 15,097 proposed blocks, 10,793 unique transactions, and 2,380,014 transaction-block pairings. We find that exclusive transactions—transactions present only in blocks proposed by a single builder — account for 85% of the fees paid by all transactions included in winning blocks. We also find that a surprisingly large number of user transactions are delayed: although proposed during a bidding cycle, they are not included in the corresponding winning block. Many such delayed transactions are exclusive to a losing builder. We also identify two arbitrage bots trading between decentralized (DEX) and centralized exchanges (CEX). By examining their bidding dynamics, we estimate that the implied price at which these bots trade USDC/WETH and USDT/WETH on CEXes is between 3.4 and 4.2 basis points better than the contemporaneous price reported on Binance.
Keywords:
Ethereum, Proposer-Builder Separation (PBS), Order Flow.
@@ 1 Introduction
inlineinlinetodo: inlinebefore submitting, we should make sure that we are consistent with the naming of builders and relaysinlineinlinetodo: inlinechange ”block value” to ”block’s potential revenues” (or somethign else)
Blockchains are often described as immutable public ledgers because transactions included in the canonical chain cannot be altered. This static characterization, however, conceals the probabilistic and competitive process that determines which transactions become immutable. In Ethereum, for each block added to the canonical chain, up to 5,000 candidate blocks are constructed, propagated, but ultimately discarded along with their transactions. Each of these discarded blocks represents a counterfactual, unrealized history. In one candidate block, a user might secure a highly profitable trade; in another, the same user could suffer a loss due to an attack. When such differences arise, the choice of winning block has significant consequences for users. Conversely, if blocks differ only in a handful of arbitrage transactions, the outcome may matter little to regular users. Understanding how proposed blocks differ and the resulting implications for users is essential to understanding a system that facilitates roughly USD 5 billion in daily financial exchanges.
In this paper, we perform the first analysis of transactions included in non-winning Ethereum blocks. We collect 15,097 non-winning blocks submitted between Ethereum blocks 21,322,622 to 21,322,660 (38 winning blocks), approximately from 2:37:35 PM to 2:45:25 PM UTC on December 3, 2024. Our dataset contains 10,793 individual transactions (identified by their hash) and 2.380.014 individual transactions/blocks combinations, which will be our main unit of analysis. During the approximately eight-minute we consider, the price of ETH against USDC on Binance showed notable volatility. It began at 3,517.45 (2:37:35 PM UTC), rose by 1.06% to 3,554.92 (2:42:12 PM), dipped by 0.37% to 3,541.68 (2:43:13 PM), and then increased again by 0.60% to close at 3,562.79 (2:45:25 PM) (see Figure 1).
Before presenting our main results, we review the process by which blocks are submitted and selected for inclusion (readers familiar with it may skip this section) and describe our dataset in greater detail, including its limitations. After presenting our findings, we discuss their relevance to the broader debate about Ethereum’s current design and their relation to the existing academic literature.
Figure 1: ETH/USDC price on Binance during our study period
1.1 Background: the journey of an Ethereum transaction
On Ethereum, the right to add a new block to the canonical chain is assigned to a proposer (or validator), randomly selected from addresses that have locked (or staked) a sufficient amount of ETH, Ethereum’s native token. The proposer can assemble the block using transactions from the public mempool, a repository of pending transactions submitted by Ethereum users. However, in over 90% of cases, the proposer delegates block creation to specialized entities called builders, who gather both user transactions and those submitted by searchers. Searchers are automated bots that generate transactions in response to market conditions (for example, price fluctuations) or in response to other pending transactions in the mempool. For builders, selecting which transactions to include and their order is a complex optimization problem: block space is limited, not all transactions can be included, and some may conflict with others.
Users and searchers interact with builders in different ways. Some users submit their transactions to the Ethereum public mempool, where builders observe them. However, this exposes them to potential attacks. [^1] For example, if a user swaps assets on a blockchain-based financial market (an Automated Market Maker, or AMM), a malicious searcher can execute a “sandwich” attack: it front-runs the user’s swap by making the same trade first, then back-runs it with the opposite swap. This allows the attacker to buy low and sell high, forcing the victim to trade at worse prices. To mitigate this risk, users may instead use private mempools, services that connect them directly with multiple builders. Searchers always submit their transactions privately, either through known private mempools or via direct communication with one or more builders. Every transaction includes either a fee or a payment to the “Fee Recipient” address, a block-level variable set by the builder (typically an Ethereum address they control). As a result, builders earn fees from users and searchers whenever they include their transactions in a block that is added to the blockchain.
The relation between builders and proposers is facilitated by relays. Their primary role is to prevent proposers from altering the submitted blocks, for example, by changing the “Fee Recipient” variable. [^2] Technically, this is achieved by requiring proposers to sign an empty block, which is then completed by the relay. For more details, see https://docs.flashbots.net/flashbots-mev-boost/relay. Once a builder creates a block, it submits it to a relay along with a proposed payment for the proposer. The relay then forwards the highest-paying block and its corresponding payment to the proposer upon request. This process resembles an ascending price auction because the relay continuously broadcasts the value of the highest-paying bid—though this broadcast is observed with a delay due to network latency. An auction cycle lasts approximately 12 seconds, after which the winning block is selected, and a new auction begins. During each auction cycle, builders continuously resubmit blocks, modifying the transactions included and bid amounts. Latency plays a crucial role, as proposers are geographically dispersed, and arbitrage transactions often depend on rapid price movements in traditional “off-chain” financial markets. To account for geographic dispersion and fast-moving arbitrage opportunities, multiple relays operate worldwide, with builders submitting to several at once. As a result, thousands of blocks are created and discarded for every block added to the Ethereum blockchain.
1.2 Our dataset
Our primary dataset includes all blocks (winning and non-winning) submitted via a relay called “Agnostic relay” during the study period. [^3] Founded in 2022, Agnostic relay held 20% of the relay market in 2023, declining to around 5% during our study period. We complement our primary dataset with several publicly available data sources:
  • Winning blocks for our study period, including those not supplied by Agnostic relay. Additionally, we track 10 winning blocks beyond our study period to determine whether transactions in our primary dataset were eventually included in the blockchain.
  • Hashes of blocks submitted to other relays (Flashbots, BloXroute, Manifold, Eden, Ultra Sound, SecureRpc, and Aestus), along with the name of the builders submitting them and their bids. Here, a block’s hash serves as its unique identifier, as two blocks that differ in a single transaction or their bids have different hashes. We can, therefore, track whether a block in our primary dataset was also submitted to other relays and its associated bid. In particular, we find that the 15,097 submitted blocks included in our primary dataset constitute 28.3% of all blocks submitted via major relays during the study period.
  • Data on transactions submitted via the public mempool and two known private mempools (MEV Blocker and Flashbots Protect), obtained from https://mempool.guru/ and https://Dune.com.
  • Second-by-second price data from Binance.
We also simulate the non-winning blocks in our dataset to assess how the execution of a transaction — particularly a swap on a blockchain-based financial market — varies depending on the block in which it is included. We simulate the block using an Ethereum Archive Node running the Erigon client, forking at the appropriate block heights, and replaying the blocks.
@@ 1.2 Our dataset
Our primary dataset has two main limitations. First, we observe no submitted blocks for 11 bidding cycles because the proposers assigned to those slots were not connected to the Agnostic relay, which therefore did not collect block submissions. Second, it contains proportionally few blocks from Beaverbuild, a major block builder during the study period: while Beaverbuild accounted for 40.2% of all blocks submitted to major relays, it contributed only 9.4% of the blocks submitted to Agnostic (and thus included in our dataset). We later discuss how latency may explain Beaverbuild’s limited presence on Agnostic. With these limitations in mind, we can now discuss our main results.
1.3 Summary of the main results.
The first observation is that the set of proposed transactions differs greatly from the subset of these transactions that appear in winning blocks. The vast majority (87.7%) of the transactions in our primary dataset that appear in winning blocks are users’ transactions, and only a few are searchers’ transactions. [^4] We classify an address as a searcher if it trades on decentralized exchanges (DEXes) using a smart contract whose source code is not disclosed on https://Etherscan.com. In other words, part of the searcher’s logic is embedded in a smart contract but hidden by making only its bytecode available (the bytecode is the compiled, machine-readable version of the smart contract). Swapping on a DEX is necessary for all common types of “searching,” such as sandwich attacks, arbitrage, backrunning, and triggering liquidations on lending protocols. Also, only few (14.4%) of users’ transactions are swaps. Instead, among the set of proposed transactions, about half are from users and the rest from searchers. Also, about half of the proposed transactions are swaps, which are highly concentrated among a few searchers: the top four addresses alone account for 3,653 swap transactions, or approximately 70% of the total swaps.
A notable feature of our dataset is that 2,981 transactions — 35% of all unique transactions — originate from just two addresses that almost exclusively perform swaps. We identify these addresses as searchers integrated with the two main builders in our dataset, Rsync and Titan, and refer to them as “Rsync-bot” and “Titan-bot,” respectively. Titan-bot sends transactions exclusively to Titan, while Rsync-bot sends transactions mainly to Rsync but also to Titan. By examining transaction logs, we find that all Rsync-bot transactions sent to Titan were also sent to Rsync. Notably, in every instance where the Rsync-bot submits the same transaction to both builders, the fee offered to Titan is always lower than the one offered to Rsync, with an average difference of 18%. We interpret this gap as the bot’s target profit margin when interacting with builders that are not integrated with it.
Our dataset allows us to distinguish between private transactions shared among multiple builders and exclusive transactions — those available to a single builder. Among the 5,576 transactions in our primary dataset that were included in a winning block, Mempoolguru classifies 4,124 as public and 1,452 as non-public. For the non-public subset, we check whether each transaction appeared in blocks proposed by multiple builders (private) or only in blocks from a single builder (exclusive) during the bidding cycle that led to their inclusion on chain. We find that 45% of non-public transactions included in winning blocks are exclusive. These transactions are responsible for 85% of the winning blocks’ revenues (i.e., the sum of the payments to the builder), while private but not exclusive transactions constitute 10% of revenues.
Interestingly, we find a number of transactions that change status over time: from exclusive to private and even from exclusive to public. Specifically, we identify 19 transactions that are exclusive to a single builder during the initial part of the auction cycle and later appear in blocks from other builders, thereby transitioning from exclusive to private within the same auction cycle. In the majority of cases, these transactions are submitted via Flashbots Protect and are initially exclusive to the Flashbots builder before being shared with others. We also observe 12 users’ transactions that appear exclusive to a single builder for one or two bidding cycles, and later appear in the public mempool when the builder does not win. We speculate that a private mempool operator shares each transaction exclusively with a single builder, falling back to the public mempool if that builder does not win. [^5] The presence of transactions that change status is noteworthy because four of the five major private mempool operators explicitly state that they share transactions with multiple builders (those are MEV Blocker, Flashbots Protect, Blink, and Merkel; for an analysis of the difference between these operators see Janicot and Vinyas, 2025). In contrast, the fifth major private mempool operator (Metamask Smart Transactions) does not publicly document its sharing policies.
One of our main findings is that approximately 20% of users’ transactions are delayed: they are proposed during a bidding cycle before the winning block is chosen, but they are not included in the corresponding winning block. Strikingly, 30% of these delayed transactions initially appear as exclusive. This is difficult to rationalize, as users do not typically trade based on superior information, and hence gain no advantage from sending transactions exclusively to a single builder rather than sharing them privately with multiple builders or broadcasting them publicly. Even more surprisingly, as already discussed, some of these exclusive transactions later appear in the public mempool. These users’ transactions are therefore delayed and also exposed to attacks.
We also examine what determines whether a swap executes in one block but not another, and the price at which it executes. Regression results show that a swap in the same direction as Rsync-bot or Titan-bot is approximately 18% less likely to execute when included in a block built by Rsync or Titan that also includes a transaction from their respective bot. In contrast, a swap in the opposite direction is 1% more likely to execute under the same conditions. A similar, albeit smaller, pattern emerges for prices: swaps in the opposite direction as the two bots trade at better prices when included in blocks by either Titan or Rsync that also include transactions of their respective bots, while the opposite is true for swaps in the same direction as the bots.
Finally, we study the competition between Rsync-bot and Titan-bot. Because each bot is integrated with its respective builder, its fee for block inclusion measures its risk-adjusted expected profit. By observing each bot’s swap volume, the on-chain execution price, and its inclusion payment, we infer an implied centralized exchange (CEX) price for the off-chain leg of the trade. This allows us to compare the bots’ implied CEX prices to the contemporaneous Binance price and examine how this difference evolves during the bidding cycle. Focusing on instances where both bots compete to rebalance either a USDT/WETH pool or a USDC/WETH pool, we find that the implied CEX price for a marginal (zero-volume) trade at second 12 of the bidding cycle is approximately 2.3 basis points better than Binance’s price for Rsync-bot and 1.5 basis points better for Titan-bot. For a 10 ETH trade, the corresponding improvements over Binance are 2 basis points for Rsync-bot and 1.2 for Titan-bot. For other token pairs, for which we have much fewer observations than for USDT/WETH and USDC/WETH, the evidence is ambiguous and we find no indication that the bots trade at a better or worse price than the one reported by Binance.
@@
Finally, we analyze the competition between Rsync-bot and Titan-bot. Since each bot is integrated with its respective builder, its payment for block inclusion equals its risk-adjusted expected profit. By observing each bot’s swap volume, on-chain execution price, and payment for inclusion, we infer an implied centralized exchange (CEX) price — i.e., the price of the off-chain leg of the trade. This inferred CEX price can then be compared to the contemporaneous Binance price during the bidding cycle. Focusing on cases where both bots compete to rebalance either a USDT/WETH or a USDC/WETH pool, we find that the implied CEX price is approximately 4.2 basis points better than Binance’s price for Rsync-bot, and 3.4 basis points better for Titan-bot.
1.4 Contribution
@@
Ethereum’s process of block creation, submission, and selection— known as Proposer-Builder Separation (PBS) — has been the subject of considerable debate and criticism. [^6] For a summary of this debade, with a particular emphasis on the economic aspects, see John et al. (2025). Our results contribute directly to these discussions.
A primary concern is the extreme concentration in the builders’ market: the top two builders (Titan and Beaverbuild) currently produce between 80% and 90% of blocks added to the Ethereum blockchain. These entities wield substantial influence over transaction inclusion, posing a threat to the blockchain’s foundational goals of decentralization and permissionlessness. This concentration is often attributed to economies of scale from exclusive transactions: the more often a builder wins, the more likely it is to attract transactions not shared with competitors, and the more this builder wins. Until now, this hypothesis has remained untested, as on-chain data alone does not allow distinguishing between transactions shared privately with multiple builders and truly exclusive ones. Our analysis fills this gap by showing that 85% of the revenues from winning blocks in our dataset, measured by total payments to builders, is generated by exclusive transactions.
@@
A second major criticism of PBS is that it enables searchers to efficiently extract value from users. Quantitatively, the main concern is arbitrageurs profiting at the expense of on-chain liquidity providers by exploiting price discrepancies between decentralized exchanges (DEXes) and centralized exchanges (CEXes). The resulting loss to liquidity providers is referred to as Loss-vs-Rebalancing (LVR), a concept introduced by Milionis et al. (2022). Subsequent studies have attempted to identify arbitrageurs (Heimbach et al., 2024) and quantify LVR (Canidio and Fritsch, 2023; Fritsch and Canidio, 2024). These analyses, however, rely on strong assumptions. This is because on-chain data alone reveals only that a swap occurred, not whether it was part of an arbitrage strategy. Furthermore, the prices at which arbitrageurs trade on centralized venues are usually unobservable. Our dataset overcomes these limitations in two ways. First, we can identify arbitrage bots by observing competition between them for the same arbitrage opportunities. [^7] In this respect, our approach resembles that of Aquilina et al. (2022), who identify high-frequency trading races in traditional financial markets by analyzing failed transactions. Second, by focusing on builder-integrated searchers engaging in CEX-DEX arbitrage, we can infer their expected profits and thus back out the effective price at which they trade on the centralized leg.
A final criticism is that a large section of the PBS supply chain is opaque and may not operate in the users’ interest. Our paper illuminates some of its aspects and shows that, indeed, there are reasons for concern regarding the handling of users’ transactions.
1.5 Additional relevant literature
@@
The dynamics of builder competition and block inclusion on the Ethereum blockchain have been studied extensively. However, to the best of our knowledge, we are the first to systematically collect data on transactions contained in non-winning blocks. Yang et al. (2024) is the only other paper that analyzes the content of non-winning blocks. They examine competition among builders by comparing builders’ bids to the blocks’ revenues. Several other papers explore different aspects of builder competition. For example, Wu et al. (2024a) and Wu et al. (2024b) study builders’ bidding strategies, both theoretically and through simulation. Titan and Frontier Research (2023) and Öz et al. (2024) show that builder profits depend on how they source transactions, using a classification approach similar to ours. Bahrani et al. (2024) analyze builder competition theoretically and derive conditions under which the builder and proposer markets become concentrated in equilibrium. While we also study builder competition, our focus is on block-building strategies, for instance, whether they are vertically integrated with searchers and the behavior of these searchers.
@@
Similarly to our analysis of the competition between Rsync-bot and Titan-bot, Capponi et al. (2024) also studies searchers competing to exploit CEX-DEX arbitrage opportunities. However, they consider an earlier period in which searchers competed by submitting transactions to the public mempool and engaging in priority fee auctions — i.e., because all transactions are included in the block and ordered by priority fee, the winning searcher is the one paying the highest fee. The current market structure and nature of the competition between searchers are very different, with searchers submitting directly to builders, builders selecting which searchers’ transactions to include, and frequent integration between searchers and builders. [^8] Also notable is Capponi et al. (2023), in which the same authors develop a theory of searchers’ competition when transactions are sent privately to builders.
@@
Also related are Öz et al. (2023) and Schwarz-Schilling et al. (2023), who study how proposers decide when to request the winning block. They argue that this choice is partially strategic, as builders’ bids tend to increase during the auction cycle. Delaying the selection of the winning block can, therefore, lead to higher payments to the proposer. Wahrstätter et al. (2023) and Heimbach et al. (2023) provide an empirical analysis of Ethereum block production from September 2022 to May 2023, a period during which proposer-builder separation (PBS) became the dominant mechanism for block construction on Ethereum.
Finally, Pai and Resnick (2024) develops a theory of integration between builders and searchers, based on the idea that an integrated searcher’s payment to a builder reflects the full value of including the transaction. In contrast, non-integrated searchers pay less than the full value to retain a profit if their transaction is included. This has implications for builder competition, as it suggests that builders integrated with searchers enjoy a competitive advantage over non-integrated builders. These insights are relevant to our analysis. In particular, we observe integrated searchers submitting the same arbitrage transaction both to the builder with whom they are integrated and to others with whom they are not. In such cases, the payment offered to the integrated builder is consistently higher than that offered to non-integrated builders. We interpret this higher payment as reflecting the true value of transaction inclusion.

### v2

@@ Abstract
We collect and study 15,097 blocks proposed for inclusion in the Ethereum blockchain during 8 minutes on December 3rd, 2024, corresponding to 39 added blocks. These proposed blocks contain 10,793 unique transactions, and 2,380,014 transaction-block pairings, our primary unit of analysis. We find that 20% of user transactions are delayed: although proposed during a bidding cycle, they are not included in the corresponding winning block. Approximately 30% of such delayed transactions are exclusive to a losing builder. We also identify two arbitrage bots trading between decentralized (DEX) and centralized exchanges (CEX), which are responsible for a significant fraction of the value of proposed blocks. By examining their bidding dynamics, we estimate that the implied price at which these bots trade USDC/WETH and USDT/WETH on CEXes is approximately 2.8 basis points better than the contemporaneous Binance price.
@@ 1 Introduction
inlineinlinetodo: inlinebefore submitting, we should make sure that we are consistent with the naming of builders and relaysinlineinlinetodo: inlinechange ”block value” to ”block’s potential revenues” (or somethign else)inlineinlinetodo: inlineMissing: more info on transaction delay. A plot, and also what these delays constitute in terms of feeinlineinlinetodo: inlineMissing: something that gives a general sense of how much variation there is in terms of prob of success and price (then we can jump to explaining this variation)
Blockchains are often described as immutable ledgers because once a transaction is included in the canonical chain it cannot be altered. This static view, however, obscures the probabilistic and competitive process that determines which transactions become immutable. In Ethereum, this process is governed by the interaction of proposers, builders, and relays. The block proposer has the right to add the next block to the chain and is randomly selected among those who have staked at least 32 ETH. Builders are specialized entities that gather transactions and use algorithms to decide which to include and in what order. Relays act as auctioneers: they collect blocks and payments offered to the proposer (bids) from builders, continuously broadcast the current highest bid, and forward the corresponding block to the proposer when requested. [^1] Relays are “trusted” because they must also prevent proposers from tampering with submitted blocks. For more details, see https://docs.flashbots.net/flashbots-mev-boost/relay. Each auction cycle lasts about 12 seconds, after which the winning block is added to the chain and the process begins again.
Within a cycle, builders continually collect new transactions from users and from searchers, automated bots that generate transactions in response to market conditions (such as price movements) or to other pending transactions. They use these new transactions to update and resubmit blocks. Latency is crucial, as proposers are globally distributed and arbitrage opportunities often depend on rapid movements in off-chain financial markets. To mitigate these frictions, multiple relays operate worldwide, and builders typically submit to several at once. As a result, for each block added to the canonical chain, up to 5,000 candidate blocks are constructed, propagated, and ultimately discarded along with their transactions. Each of these discarded blocks represents an alternative, unrealized history. Measuring how much these unrealized histories differ from the one that ultimately appears on-chain is a key research question for a system that supports about USD 5 billion in daily financial exchanges. Yet, until now, data limitations have prevented researchers from addressing it.
@@
In this paper, we overcome this limitation by assembling a novel dataset of proposed blocks submitted between Ethereum blocks 21,322,622 and 21,322,660 (39 winning blocks), roughly from 2:37:35 PM to 2:45:25 PM UTC on December 3, 2024. This dataset contains 15,097 blocks, 10,793 unique transactions (identified by their hash) and 2,380,014 transaction–block pairs, which constitute our main unit of analysis. Crucially, the blocks in our dataset are not a random sample: they are all blocks submitted to a relay called Agnostic. [^2] Founded in 2022, Agnostic relay held about 20% of the relay market in 2023, declining to around 5% during our study period. This introduces two limitations. First, we observe no submitted blocks for 11 bidding cycles because the proposers assigned to those slots were not connected to Agnostic, which therefore did not collect block submissions. Second, it contains proportionally few blocks from Beaverbuild, a major builder during the study period: in bidding cycles for which we have data, Beaverbuild accounted for 40.2% of all blocks submitted to major relays but it contributed only 9.4% of the blocks submitted to Agnostic (and thus included in our dataset). We later discuss how latency may explain Beaverbuild’s limited presence on Agnostic. With these limitations in mind, we can now discuss our main results.
Our main research question is whether the process of block selection matters for users. We find that it does. Approximately 20% of users’ transactions are delayed: they first appear during an auction cycle in some proposed blocks but not in the winning one, delaying their inclusion on-chain by one or more bidding cycles. Strikingly, about 30% of these delayed transactions initially appear as exclusive, i.e., they are present only in blocks submitted by a single builder. We also examine what determines whether a user’s swap executes in some blocks but not others, and at what price it executes. We document substantial variance in both execution success and prices across blocks containing the same transaction.11todo: 1add values Regression results show that part of this variance can be explained by what other transaction is included in a block: swaps in the same direction as the two main arbitrage bots in our data (see the next paragraph) are about 18% less likely to execute when included in a block that also contains transactions from those bots. By contrast, swaps in the opposite direction are about 1% more likely to execute under the same conditions. A similar, though smaller, pattern holds for prices: swaps in the opposite direction of the bots trade at better prices when included in blocks that also contain their transactions, while swaps in the same direction trade at worse prices. Hence, the process of block selection materially affects users, particularly through the timing of transaction execution.
We then turn to the determinants of which block wins. To do so, we exploit a notable feature of our dataset: 2,981 transactions—35% of all unique transactions—originate from just two addresses that almost exclusively perform swaps and together account for roughly 60% of all transaction fees in the blocks where they appear. The first address, which we label “Titan-bot,” submits transactions only to the builder Titan. The second, “Rsync-bot,” submits transactions almost exclusively to a builder called Rsync, although a few of its transactions occasionally appear in blocks built by both Rsync and Titan. In those cases, however, the fee offered to Titan is always lower than the fee offered to Rsync, with an average difference of 18%. [^3] It is common knowledge that Rsync owns and operates several trading bots. Titan, by contrast, claims not to own or operate any bot. Wu et al. (2025) also identify Titan-bot (“Kyle” in their labels) as exclusive to Titan.
@@
Since these two bots generate most of the fees in each block, their competition sheds light on which builder wins the auction. We assume that each bot’s payment to its builder for block inclusion equals its risk-adjusted expected profit. [^4] This assumption is justified if the bot is owned by the builder itself (Pai and Resnick, 2024), or if the builder rebates to the bot an appropriate share of the fee. For example, a rebate may be calculated in a second-price auction style by comparing the block’s value with and without the bot’s transaction. This is the approach taken by Buildernet, a recent open-source builder operating inside a Trusted Execution Environment (see https://buildernet.org/docs/refunds). By observing each bot’s swap volume, on-chain execution price, and payment for inclusion, we infer an implied centralized exchange (CEX) price—that is, the price of the off-chain leg of the arbitrage trade. We then compare this implied price to the contemporaneous Binance price during the bidding cycle. We find that when both bots compete to rebalance either a USDT/WETH or a USDC/WETH pool, the implied CEX price is roughly 2.8 basis points more favorable than Binance’s contemporaneous price for both bots.
1.1 Relevant literature
@@
Ethereum’s process of block creation, submission, and selection— known as Proposer-Builder Separation (PBS) — has been the subject of considerable debate (for a summary of this debate, with a particular emphasis on the economic aspects, see John et al. (2025)). Our results contribute directly to these discussions.
@@
One of the main criticism of PBS is that it enables searchers to efficiently extract value from users. Quantitatively, the main concern is arbitrageurs profiting at the expense of on-chain liquidity providers by exploiting price discrepancies between decentralized exchanges (DEXes) and centralized exchanges (CEXes). The resulting loss to liquidity providers is referred to as Loss-vs-Rebalancing (LVR), a concept introduced by Milionis et al. (2022). Subsequent studies have attempted to identify arbitrageurs (Heimbach et al., 2024, Wu et al. (2025)) and quantify LVR (Canidio and Fritsch, 2023; Fritsch and Canidio, 2024). These analyses, however, rely on strong assumptions. This is because on-chain data alone reveals only that a swap occurred, not whether it was part of an arbitrage strategy. Furthermore, the prices at which arbitrageurs trade on centralized venues are usually unobservable. Our dataset overcomes these limitations in two ways. First, we can identify arbitrage bots by observing competition between them for the same arbitrage opportunities. In this respect, our approach resembles that of Aquilina et al. (2022), who identify high-frequency trading races in traditional financial markets by analyzing failed transactions. Their rationale is that when multiple arbitrageurs try to exploit the same arbitrage opportunity, only one transaction will succeed, and all others will fail. Similarly, we observe multiple arbitrage bots trying to exploit the same opportunity by submitting transactions, with only one of them eventually successfully included on-chain. Second, by focusing on builder-integrated searchers engaging in CEX-DEX arbitrage, we can infer their expected profits and thus the price at which they trade on the centralized leg.
@@
The dynamics of builder competition and block inclusion on the Ethereum blockchain have been studied extensively. However, to the best of our knowledge, we are the first to systematically collect data on transactions contained in non-winning blocks. Yang et al. (2024) is the only other paper that analyzes the content of non-winning blocks. They examine competition among builders by comparing builders’ bids to the blocks’ revenues. Several other papers explore different aspects of builder competition. For example, Wu et al. (2024a) and Wu et al. (2024b) study builders’ bidding strategies, both theoretically and through simulation. Bahrani et al. (2024) analyze builder competition theoretically and derive conditions under which the builder and proposer markets become concentrated in equilibrium.
@@
Similarly to our analysis of the competition between Rsync-bot and Titan-bot, Capponi et al. (2024) also studies searchers competing to exploit CEX-DEX arbitrage opportunities. However, they consider an earlier period in which searchers competed by submitting transactions to the public mempool and engaging in priority fee auctions — i.e., because all transactions are included in the block and ordered by priority fee, the winning searcher is the one paying the highest fee. The current market structure and nature of the competition between searchers are very different, with searchers submitting directly to builders, builders selecting which searchers’ transactions to include, and frequent integration between searchers and builders. [^5] See also Capponi et al. (2023), who develop a theory of searchers’ competition when transactions are sent privately to builders.
@@
Also related are Öz et al. (2023) and Schwarz-Schilling et al. (2023), who study how proposers decide when to request the winning block. They argue that this choice is partially strategic, as builders’ bids tend to increase during the auction cycle. Delaying the selection of the winning block can, therefore, lead to higher payments to the proposer. Wahrstätter et al. (2023) and Heimbach et al. (2023) provide an empirical analysis of Ethereum block production from September 2022 to May 2023, a period during which proposer-builder separation (PBS) became the dominant mechanism for block construction on Ethereum. Finally, among the studies dealing with the relationship between searchers and builders, the most relevant ones are Pai and Resnick (2024) and Wu et al. (2025). Pai and Resnick (2024) develops a theory of integration between builders and searchers based on the idea that an integrated searcher’s payment to a builder reflects the full value of including the transaction. Wu et al. (2025) uses on-chain data to identify searchers performing CEX-DEX arbitrage, and study their relationship with the major builders.
Finally, in our companion paper Pahari and Canidio (2025), we use the same data to measure the share of transactions that are exclusive to a builder, meaning they appeared only in blocks proposed by the winning builder during the auction cycle. This question is particularly important given the high concentration in the builders’ market: the top builder (Titan) produces about 50% of all blocks, while the top three account for more than 90%. [^6] For up-to-date statistics, see https://www.relayscan.io/. Buildernet is a joint project by Flashbots (who operated an independent builder until mid-2025) and Beaverbuild (who both participate in Buildernet and run their own builder). Such concentration gives a few entities disproportionate influence over which transactions are included, undermining Ethereum’s goals of decentralization and permissionlessness. We find that between 77% and 84% of the total value of transactions included on-chain is attributable to exclusive transactions, confirming that exclusivity is the main driver of builders’ profits and of PBS auction outcomes.